OpenStudy (anonymous):

Given the linear equation x + y = 5, find the slope of its graph.

myininaya (myininaya):

Do you know that y=mx+b is called slope-intercept form? It is because m tells us the slope . And b tells us about the y-intercept.

myininaya (myininaya):

Can you try to write x+y=5 in the form y=mx+b ?

OpenStudy (anonymous):

xy=5

myininaya (myininaya):

addition doesn't turn into multiplication for example 3+2 is not the same as 3(2) you know since 3+2=5 and 3(2)=6

myininaya (myininaya):

Your objective is to solve x+y=5 for y. Try either subtracting/adding something on both sides to do this.

OpenStudy (anonymous):

its just 5?

myininaya (myininaya):

Not sure what you are talking about? x+y is 5 yes but we want to write x+y=5 in the form y=mx+b that means basically I need you to attempt to solve x+y=5 for y. The hint I gave is for you to pick something to subtract on both sides.

myininaya (myininaya):

do you see the addition of x on the same side as y? Try subtracting x on both sides.

myininaya (myininaya):

@Alexisboo1998 ?

OpenStudy (anonymous):

i just picked 5 idk